### v2

# Convolutional and Deep Learning based techniques for Time Series Ordinal Classification

Abstract
Time Series Classification (TSC) covers the supervised learning problem where input data is provided in the form of series of values observed through repeated measurements over time, and whose objective is to predict the category to which they belong. When the class values are ordinal, classifiers that take this into account can perform better than nominal classifiers. Time Series Ordinal Classification (TSOC) is the field covering this gap, yet unexplored in the literature. There are a wide range of time series problems showing an ordered label structure, and TSC techniques that ignore the order relationship discard useful information. Hence, this paper presents a first benchmarking of TSOC methodologies, exploiting the ordering of the target labels to boost the performance of current TSC state-of-the-art. Both convolutional- and deep learning-based methodologies (among the best performing alternatives for nominal TSC) are adapted for TSOC. For the experiments, a selection of $29$ ordinal problems from two well-known archives has been made. In this way, this paper contributes to the establishment of the state-of-the-art in TSOC. The results obtained by ordinal versions are found to be significantly better than current nominal TSC techniques in terms of ordinal performance metrics, outlining the importance of considering the ordering of the labels when dealing with this kind of problems.
Index Terms:
time series machine learning, time series analysis, time series classification, ordinal classification
Appendix A Data description
First of all, we made a dataset selection from two time series repositories, the UEA/UCR TSC archive [^1] https://timeseriesclassification.com, and the Monash/UEA/UCR TSER archive [^2] http://tseregression.org/. For the TSC repository, we analysed the output variable of each problem to select those with an ordinal nature. We have identified seven univariate, and two multivariate datasets. Moreover, we have chosen 13 additional datasets from the TSER archive. Given that it is a regression repository, we have discretised the original continuous output variable, grouping it into five uniformly distributed bins. Even though there are more regression datasets, we could only select these 13 as it is difficult to achieve a reasonable class distribution for the rest (the output values are too close to obtain more than three representative classes). Furthermore, the datasets with unequal length time series and missing values are also discarded. Other five TSOC problems built from historical price data from five of the most important companies in the stock market are considered: Apple (AAPL), Amazon (AMZN), Google (GOOG), META and Microsoft (MSFT), and other two were built from data collected from the National Data Buoy Center (NDBC).
Data for the five TSOC problems built from historical price data has been collected from Yahoo Finance [^3] https://es.finance.yahoo.com/ website, extracting weekly price data from each stock earliest available date to March 2023. The time series of each stock data are formed by its price returns over 53 weeks (around one year) before a varying date $t$. The output ordinal label corresponds to the price return in $t$. This value is discretised according to a set of predefined symmetrical thresholds $(-\infty,-0.05,-0.02,0.02,0.05,\infty)$. Finally, two more TSOC problems were built from data collected from the National Data Buoy Center (NDBC) [^4] https://www.ndbc.noaa.gov/. Input multivariate time series are seven climatological variables collected from a reanalysis model, as detailed in [gomez2022simultaneous]. Some of these variables are the air temperature, the pressure, or the relative humidity. Each time series comprises four values per day during four weeks, i.e. the length of the time series is $112$. The output label indicates the energy fluctuation level during the data collection period, with 00 being the minimum and $3$ the highest level. The second problem also comprises multivariate time series. Same seven climatological variables are considered for this problem. This time, the length of the time series is $28$, as collects values during only one week (four values per day). The purpose is to estimate the wave height level during that period of time, from 00 (the lowest height) to $3$ (the highest height). Note that the outputs of both datasets have been discretised in bins with identical widths.
Table I presents the datasets information used in this study.
Dataset #Train #Test $Q$ $T$ $C$
TSC AtrialFibrillation 15 15 3 640 2
DistalPhalanxOAG 400 139 3 80 1
DistalPhalanxTW 400 139 6 80 1
EthanolConcentration 261 263 4 1751 3
EthanolLevel 504 500 4 1751 1
MiddlePhalanxOAG 400 154 3 80 1
MiddlePhalanxTW 399 154 6 80 1
ProximalPhalanxOAG 400 205 3 80 1
ProximalPhalanxTW 400 205 6 80 1
TSER AcousticContaminationMadrid 166 72 5 365 1
AppliancesEnergy 95 42 5 144 24
BarCrawl6min 140 61 5 360 3
Covid3Month 140 61 5 84 1
DhakaHourlyAirQuality 1447 621 5 24 1
GasSensorArrayAcetone 324 140 3 7500 1
GasSensorArrayEthanol 324 140 3 7500 1
NaturalGasPricesSentiment 65 28 5 20 1
ParkingBirmingham 1391 597 5 14 1
SolarRadiationAndalusia 470 202 5 365 2
TetuanEnergyConsumption 254 110 5 144 5
WaveDataTension 1325 568 5 57 1
WindTurbinePower 596 256 5 144 1
new TSOC AAPL 1720 431 5 53 1
AMZN 1035 259 5 53 1
GOOG 732 183 5 53 1
META 408 103 5 53 1
MSFT 1501 376 5 53 1
USASouthwestEnergyFlux 327 141 4 112 7
USASouthwestSWH 1310 562 4 28 7
Table I: Characteristics of the datasets used in this work. #Train stands for the number of training patterns, #Test that of the test set, $Q$ the number of classes, $T$ for the time series length, and $C$ is the number of channels.
Appendix B TSOC methods comparison tables
To further analyse the performance of TSOC approaches based on the main characteristics of the datasets, we have divided the performance results to gain insights into the characteristics under which different algorithms perform well. These analyses have been conducted using the average rank in terms of MAE. Table II shows the differences when grouping by time series length. As observed, O-MiniROCKET performs better with shorter series (length $<$ 1000), while O-MultiROCKET is better suited for longer time series. Table III presents the average rank when the datasets are grouped by the number of classes. Overall, O-MiniROCKET is the best approach in all groups except for the 4-class datasets, where O-LITETime performs better. This is likely because this deep learner performs particularly well on the Ethanol datasets. Table IV includes the analysis when grouping by the number of instances in the training sets. Here, O-MiniROCKET always outperforms others with a smaller number of instances. O-InceptionTime is consistently the second best with larger training set sizes. Finally, Table V compares the ranks when the datasets are grouped by the number of channels. In this case, O-MiniROCKET also achieves the best for all the categories.
Method 0–75 (10) 75–150 (11) 150–1000 (4) $>$1000 (4)
O-ROCKET $\mathit{3.300}$ $\mathit{2.727}$ $\mathit{3.250}$ $3.500$
O-MiniROCKET $\mathbf{2.000}$ $\mathbf{1.727}$ $\mathbf{2.000}$ $\mathit{3.000}$
O-MultiROCKET $4.000$ $4.818$ $3.750$ $\mathbf{2.250}$
O-ResNet $6.100$ $5.818$ $6.500$ $6.750$
O-InceptionTime $3.400$ $4.182$ $3.500$ $3.500$
O-LITETime $4.200$ $4.818$ $5.000$ $3.750$
O-CNN $5.000$ $3.909$ $4.000$ $5.250$
Table II: Average rank MAE split by time series length.
Method 3 (6) 4 (4) 5 (16) 6 (3)
O-ROCKET $\mathit{2.500}$ $3.750$ $\mathit{3.438}$ $\mathit{1.667}$
O-MiniROCKET $\mathbf{1.667}$ $\mathit{2.750}$ $\mathbf{2.125}$ $\mathbf{1.333}$
O-MultiROCKET $4.000$ $3.750$ $3.875$ $5.333$
O-ResNet $6.000$ $6.500$ $6.125$ $6.000$
O-InceptionTime $4.333$ $3.000$ $3.562$ $4.333$
O-LITETime $5.667$ $\mathbf{2.500}$ $4.438$ $5.000$
O-CNN $3.833$ $5.750$ $4.438$ $4.333$
Table III: Average rank MAE split by number of classes.
Method 0–200 (6) 201–400 (11) 401–800 (5) $>$800 (7)
O-ROCKET $\mathit{3.000}$ $\mathit{2.636}$ $3.800$ $3.429$
O-MiniROCKET $\mathbf{2.167}$ $\mathbf{1.818}$ $\mathbf{2.200}$ $\mathbf{2.143}$
O-MultiROCKET $\mathit{3.000}$ $4.455$ $4.600$ $3.857$
O-ResNet $6.500$ $5.909$ $6.800$ $5.714$
O-InceptionTime $4.833$ $3.818$ $\mathit{2.800}$ $\mathit{3.286}$
O-LITETime $5.333$ $4.636$ $3.800$ $4.000$
O-CNN $3.167$ $4.727$ $4.000$ $5.571$
Table IV: Average rank MAE split by number of train patterns.
Method 1 (21) 2–4 (4) $>$4 (4)
O-ROCKET $\mathit{3.095}$ $\mathit{2.750}$ $3.500$
O-MiniROCKET $\mathbf{1.952}$ $\mathbf{2.500}$ $\mathbf{2.000}$
O-MultiROCKET $4.143$ $3.000$ $4.500$
O-ResNet $6.048$ $6.500$ $6.250$
O-InceptionTime $3.714$ $3.750$ $3.750$
O-LITETime $4.619$ $5.500$ $\mathit{2.750}$
O-CNN $4.429$ $4.000$ $5.250$
Table V: Average rank MAE split by number of channels.